OpenStudy (anonymous):

How would I factor this: 16(x^2+1)^2-4(2x)^2

OpenStudy (jhannybean):

\[\large 16(x^2+1)^2 -4(2x)^2\]Expand \(\large 16\color{blue}{(x^2+1)^2}-4\color{blue}{(2x)^2}\) What do you get?

OpenStudy (anonymous):

16(x^2+1)(x^2+1)-4(2x)(2x)

OpenStudy (jhannybean):

Yes, that's correct, or you can apply the formula \(\large a^2 +2ab+b^2\) to \(\large x^2+1\) It will result in \[\large 16\color{blue}{(x^4 +2x^2 +1)}-4\color{blue}{(4x^2)}\]

OpenStudy (anonymous):

I have never seen that formula before. can we solve with out it?

OpenStudy (raden):

16(x^2+1)^2 - 4(2x)^2 = 4^2(x^2+1)^2 - 2^2(2x)^2 = (4(x^2+1))^2 - (2(2x))^2 = (4x^2+4)^2 - (4x)^2 then use the difference square formula : a^^2 - b^2 = (a+b)(a-b) continue it ...

OpenStudy (raden):

* a^2

OpenStudy (jhannybean):

I guess whichever method you prefer....

OpenStudy (jhannybean):

Still there?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

still trying to figure it out. I think Im suppose to use the difference of squares thing

OpenStudy (jhannybean):

You can use any method you like, really. It'll all result in the same solution.

OpenStudy (anonymous):

Dont understand why 4^2 and 2^2 are in line 2 Dont I take the gcf out

OpenStudy (anonymous):

in Raden

OpenStudy (jhannybean):

To make it to the same power, so the terms can be combined.

OpenStudy (jhannybean):

\[\large 16(x^2+1)^2 - 4(2x)^2\] because \(\large (x^2+1)^2\) is to the second power, we can change \(\large 16 \iff 4^2\) so these two terms, \(\large (x^2+1)^2\) and \(\large 4^2\) are to the same power. Therefore they can be combined as so. \[\large 4^2(x^2+1)^2 = (4(x^2+1))^2\]

OpenStudy (jhannybean):

Same thing with \(\large 4(2x)^2\) $ can be rewritten : \(\large 4 \iff 2^2 \) So we would have: \(\large 2^2(2x)^2 \iff (2(2x))^2\)

OpenStudy (jhannybean):

It*

OpenStudy (anonymous):

Ok how do I then find the factors. I understand everything p to the last line in rd example

OpenStudy (anonymous):

up

OpenStudy (jhannybean):

So now you want to utilize the formula for difference of squares, \(\large a^2 -b^2\). You've got: \(\large [4(x^2+1)]^2- [2(2x)]^2\) First expand whats inside the brackets. \(\large (4x^2+4)^2 -(4x^2)\) this is your difference of squares equation. \(\large a^2 = (4x^2+4)^2 \ \therefore \ a = 4x^2+4\) \(\large b^2 = (4x)^2 \ \therefore \ b = 4x\)

OpenStudy (anonymous):

the answer in the back said 16(x^2-x+1)(x^2+x+1) I dont see how this works out to be hat

OpenStudy (anonymous):

not sure how they came up with that answer

OpenStudy (jhannybean):

\[\large (a+b)(a-b) = (4x^2+4+4x)(4x^2+4-4x)\]

OpenStudy (jhannybean):

Now use the foil method or grouping method to solve.

OpenStudy (jhannybean):

And easy method would be to factor out the GCF, 4.

OpenStudy (anonymous):

ok cool

OpenStudy (jhannybean):

What do you get?

OpenStudy (anonymous):

4(x^2+1+x)x^2+1-x)

OpenStudy (anonymous):

forgot the bracket in front of the x

OpenStudy (jhannybean):

You made a small mistake there. We would be factoring out a 4 from each of the two terms, \(\large (4x^2+4+4x)\) and \(\large (4x^2+4-4x)\)

OpenStudy (jhannybean):

\[\large 4(x^2+x+1) \cdot 4(x^2 -x+1) = (4 \cdot 4)(x^2+x+1)(x^2-x+1) \]

OpenStudy (anonymous):

ohh I see

OpenStudy (anonymous):

This is a great example I can use a reference. Thanks for your time. How do I save this response ?

OpenStudy (jhannybean):

are you using chrome web browser?

OpenStudy (anonymous):

safari

OpenStudy (anonymous):

I got link sent to me in the mail email

OpenStudy (anonymous):

Thanks again

OpenStudy (jhannybean):

Oh ok :)